# Interpretable and flexible non-intrusive reduced-order models using reproducing kernel Hilbert spaces

### v1

@@ 2.3.2 Known structure: feature map kernels
where ${\boldsymbol{\phi}}:{}^{n_{x}}\to{}^{n_{\phi}}$ is called the feature map and $\mathbf{G}\in{}^{n_{\phi}\times n_{\phi}}$ is a symmetric positive definite weighting matrix. It can be easily verified that feature map kernels are positive definite kernels (see, e.g., [49]). A feature map kernel results in a kernel interpolant of the form
$\mathbf{s}_{\mathbf{v}}^{\gamma}(\mathbf{x})$ $={\boldsymbol{\Omega}}^{\mathsf{T}}K(\mathbf{X},\mathbf{x})={\boldsymbol{\Omega}}^{\mathsf{T}}{\boldsymbol{\phi}}(\mathbf{X})^{\mathsf{T}}\mathbf{G}{\boldsymbol{\phi}}(\mathbf{x})=\underbrace{{\boldsymbol{\Omega}}^{\mathsf{T}}{\boldsymbol{\phi}}(\mathbf{X})^{\mathsf{T}}\mathbf{G}}_{\mathbf{C}}{\boldsymbol{\phi}}(\mathbf{x})=\mathbf{C}{\boldsymbol{\phi}}(\mathbf{x}),$ (2.12b)
where ${\boldsymbol{\phi}}(\mathbf{X})\coloneqq[~{}{\boldsymbol{\phi}}(\mathbf{x}_{1})~{}~{}\cdots~{}~{}{\boldsymbol{\phi}}(\mathbf{x}_{m})~{}]\in{}^{n_{\phi}\times m}$. Importantly, the matrix $\mathbf{C}\in{}^{n_{y}\times n_{\phi}}$ can be computed once and reused repeatedly for online kernel evaluations. After constructing $\mathbf{C}$, the cost of evaluating a feature map kernel interpolant is therefore $\mathcal{O}(n_{\phi}n_{y})$, plus the expense of evaluating ${\boldsymbol{\phi}}$ once.
@@
${\boldsymbol{\phi}}(\mathbf{x})=\begin{bmatrix}\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\end{bmatrix}\in{}^{n_{x}+n_{x}^{2}},$ (2.13)
@@
$\mathbf{s}_{\mathbf{v}}^{\gamma}(\mathbf{x})=\mathbf{C}_{1}\mathbf{x}+\mathbf{C}_{2}[\mathbf{x}\otimes\mathbf{x}],\qquad\mathbf{C}=[~{}\mathbf{C}_{1}~{}~{}\mathbf{C}_{2}~{}]\in{}^{n_{y}\times(n_{x}+n_{x}^{2})}.$ (2.14)
@@ 2.3.3 Hybrid approach
where $c_{\phi},c_{\psi}\in{}_{>0}$ are positive weighting coefficients and ${\boldsymbol{\phi}}$ is chosen to keep $n_{\phi}$ from being too large. The resulting kernel interpolant then has the form
@@
where $\mathbf{A}_{1}\in{}^{n_{x}\times n_{x}}$, $\mathbf{A}_{2}\in{}^{n_{x}\times n_{x}^{2}}$, $\mathbf{A}_{3}\in{}^{n_{x}\times n_{x}^{3}}$, and $\mathbf{A}_{4}\in{}^{n_{x}\times n_{x}^{4}}$. One option is to fully capture the structure using a quartic feature map,
${\boldsymbol{\phi}}(\mathbf{x})=\begin{bmatrix}\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\otimes\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\otimes\mathbf{x}\otimes\mathbf{x}\end{bmatrix}\in{}^{n_{x}+n_{x}^{2}+n_{x}^{3}+n_{x}^{4}}.$ (2.17)
@@
In some cases, in particular when using high-order polynomial feature maps, the kernel matrix $K(\mathbf{X},\mathbf{X})$ used for determining ${\boldsymbol{\Omega}}$ may be poorly conditioned. Increasing the regularization constant $\gamma$ can improve the conditioning of the system eq. 2.9b, but this can also degrade the accuracy of the resulting kernel interpolant. Applying a normalization to the inputs can help remedy the situation: for any injective ${\boldsymbol{\nu}}:{}^{n_{x}}\to{}^{n_{x}}$, if $K$ is positive definite, then the function $K_{\boldsymbol{\nu}}:{}^{n_{x}}\times{}^{n_{x}}\to\real$ defined by
@@
is also a positive-definite kernel function [49], and choosing ${\boldsymbol{\nu}}$ judiciously can improve the conditioning of $K_{\boldsymbol{\nu}}(\mathbf{X},\mathbf{X})$ compared to $K(\mathbf{X},\mathbf{X})$. A common choice is ${\boldsymbol{\nu}}(\mathbf{x})={\boldsymbol{\Sigma}}^{-1}(\mathbf{x}-\bar{\mathbf{x}})$, where ${\boldsymbol{\Sigma}}=\operatorname{diag}({\boldsymbol{\sigma}})\in{}^{n_{x}\times n_{x}}$ and $\bar{\mathbf{x}}\in{}^{n_{x}}$ with components
@@ 3.1 Generic projection-based reduced-order models
$\frac{\textrm{d}}{\textrm{d}t}\mathbf{q}(t)=\mathbf{f}(\mathbf{q}(t)),\qquad\mathbf{q}(0)=\mathbf{q}_{0}({\boldsymbol{\mu}}),$ (3.1)
where $\mathbf{q}:[0,T]\to{}^{n_{q}}$ is the state, $\mathbf{f}:{}^{n_{q}}\to{}^{n_{q}}$ governs the state evolution, $\mathbf{q}_{0}({\boldsymbol{\mu}})\in{}^{n_{q}}$ is the initial condition parameterized by ${\boldsymbol{\mu}}\in{\cal D}\subset{}^{n_{\mu}}$, and $T>0$ is the final desired simulation time. Models of this form often arise from semi-discretizations of time-dependent partial differential equations (PDEs), in which case the large state dimension $n_{q}$ corresponds to the fidelity of the underlying mesh. We call eq. 3.1 the full-order model (FOM).
@@
where $\mathbf{g}:{}^{r}\to{}^{n_{q}}$ and $\tilde{\mathbf{q}}:[0,T]\to{}^{r}$ is the reduced-order state, with $r\ll n_{q}$. The function $\mathbf{g}$ represents a decompression operation, mapping from reduced coordinates to the original high-dimensional space. We assume the existence of a corresponding compression map $\mathbf{h}:{}^{n_{q}}\to{}^{r}$, mapping high-dimensional states to reduced coordinates, such that $\mathbf{h}\circ\mathbf{g}$ is the identity. Importantly, $(\mathbf{g}\circ\mathbf{h})^{2}=\mathbf{g}\circ(\mathbf{h}\circ\mathbf{g})\circ\mathbf{h}=\mathbf{g}\circ\mathbf{h}$, i.e., $\mathbf{g}\circ\mathbf{h}$ is a projection. The evolution for the reduced state $\tilde{\mathbf{q}}(t)$ is then given by
@@
in which $\mathbf{h}^{\prime}:{}^{n_{q}}\to{}^{n_{q}\times n_{q}}$ is the Jacobian of $\mathbf{h}$ and where the final step comes from inserting the approximation eq. 3.2 into the FOM eq. 3.1. The resulting system
@@ 3.2 Linear and quadratic dimension reduction
where $\bar{\mathbf{q}}\in{}^{n_{q}}$ is a fixed reference vector, $\mathbf{V}\in{}^{n_{q}\times r}$ has orthonormal columns, and $\mathbf{W}\in{}^{n_{q}\times r^{2}}$ satisfies $\mathbf{V}^{\mathsf{T}}\mathbf{W}=\bf 0$. This approximation defines an $r$-dimensional quadratic manifold embedded in ${}^{n_{q}}$. An appropriate compression map corresponding to eq. 3.5 is given by
@@
a system of $r\ll n_{q}$ ODEs defined by the function $\tilde{\mathbf{f}}:{}^{r}\to{}^{r}$.
@@
$\mathbf{q}_{k}^{(\ell)}\coloneqq\mathbf{q}(t_{k};{\boldsymbol{\mu}}_{\ell})\in{}^{n_{q}},\qquad\begin{aligned} \ell&=1,\ldots,M,\\ k&=0,1,\ldots,n_{t},\end{aligned}$ (3.9)
@@
First, if $\mathbf{W}=\bf 0$, the manifold defined by $\mathbf{g}$ has no curvature and reduces to an affine subspace (or a linear subspace if $\bar{\mathbf{q}}=\bf 0$) of ${}^{n_{q}}$. In this case, we select $\mathbf{V}$ using proper orthogonal decomposition (POD) [9, 21, 56]. Define
$\mathbf{Q}\coloneqq\begin{bmatrix}(\mathbf{q}_{0}^{(1)}-\bar{\mathbf{q}})&\cdots&(\mathbf{q}_{n_{t}}^{(1)}-\bar{\mathbf{q}})&(\mathbf{q}_{0}^{(2)}-\bar{\mathbf{q}})&\cdots&(\mathbf{q}_{n_{t}}^{(M)}-\bar{\mathbf{q}})\end{bmatrix}\in{}^{n_{q}\times M(n_{t}+1)},$ (3.11)
@@
The product $\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}$ contains redundant terms, i.e., $\tilde{q}_{i}\tilde{q}_{j}$ appears twice for each $i\neq j$, which means two columns of $\mathbf{W}\in{}^{r\times r^{2}}$ act on the same quadratic state interaction in the product $\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}]$. As a consequence, the learning problem eq. 3.12 has infinitely many solutions. In practice, this issue is avoided by replacing $\otimes$ in eq. 3.5 with a compressed Kronecker product $\tilde{\otimes}$, defined by
$\tilde{\mathbf{q}}\,\tilde{\otimes}\,\tilde{\mathbf{q}}\coloneqq\left[~{}\tilde{q}_{1}^{2}~{}~{}\tilde{q}_{1}\tilde{q}_{2}~{}~{}\tilde{q}_{2}^{2}~{}~{}\dots~{}~{}\tilde{q}_{r-1}\tilde{q}_{r}~{}~{}\tilde{q}_{r}^{2}~{}\right]^{\mathsf{T}}\in{}^{r(r+1)/2},$ (3.13)
which leads to a matrix $\tilde{\mathbf{W}}\in{}^{r\times r(r+1)/2}$ such that $\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}]=\tilde{\mathbf{W}}[\tilde{\mathbf{q}}\,\tilde{\otimes}\,\tilde{\mathbf{q}}]$ for all $\tilde{\mathbf{q}}\in{}^{r}$. Then, if $\odot$ applies $\tilde{\otimes}$ column-wise, the optimization eq. 3.12 has a unique solution. Similar adjustments can be made for higher-order Kronecker products.
@@ 3.3 Intrusive reduced-order models for quadratic systems
where $\mathbf{A}\in{}^{n_{q}\times n_{q}}$, $\mathbf{H}\in{}^{n_{q}\times n_{q}^{2}}$. It is assumed that $\mathbf{H}$ is symmetric in the sense that $\mathbf{H}[\mathbf{q}\otimes\mathbf{p}]=\mathbf{H}[\mathbf{p}\otimes\mathbf{q}]$ for all $\mathbf{q},\mathbf{p}\in{}^{n_{q}}$. Models with quadratic structure arise from quadratic PDEs, but can also result from applying lifting transformations to models with other structure [30, 43]. With a linear state approximation ($\bar{\mathbf{q}}=\bf 0$ and $\mathbf{W}=\bf 0$), the ROM eq. 3.8 can be written as
@@
in which $\tilde{\mathbf{A}}=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}\in{}^{r\times r}$ and $\tilde{\mathbf{H}}=\mathbf{V}^{\mathsf{T}}\mathbf{H}[\mathbf{V}\otimes\mathbf{V}]\in{}^{r\times r^{2}}$. Constructing eq. 3.15 is an intrusive process because $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{H}}$ depend explicitly on $\mathbf{A}$ and $\mathbf{H}$; however, we need not have access to $\mathbf{A}$ and $\mathbf{H}$ to observe that the quadratic structure is preserved.
@@
$\frac{\textrm{d}}{\textrm{d}t}\tilde{\mathbf{q}}(t)&=\tilde{\mathbf{A}}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{H}}_{2}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{3}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{4}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)],\\ \tilde{\mathbf{q}}(0)&=\mathbf{V}^{\mathsf{T}}\mathbf{q}_{0}({\boldsymbol{\mu}}),\end{aligned}$ (3.16)

### v2

@@ 2.3.2 Known structure: feature map kernels
where ${\boldsymbol{\phi}}:\real^{n_{x}}\to\real^{n_{\phi}}$ is called the feature map and $\mathbf{G}\in\real^{n_{\phi}\times n_{\phi}}$ is a symmetric positive definite weighting matrix. It can be easily verified that feature map kernels are positive definite kernels (see, e.g., [50]). A feature map kernel results in a kernel interpolant of the form
$\mathbf{s}_{\mathbf{v}}^{\gamma}(\mathbf{x})$ $={\boldsymbol{\Omega}}^{\mathsf{T}}K(\mathbf{X},\mathbf{x})=\underbrace{{\boldsymbol{\Omega}}^{\mathsf{T}}{\boldsymbol{\phi}}(\mathbf{X})^{\mathsf{T}}\mathbf{G}}_{\mathbf{C}}{\boldsymbol{\phi}}(\mathbf{x})=\mathbf{C}{\boldsymbol{\phi}}(\mathbf{x}),$ (2.12b)
where ${\boldsymbol{\phi}}(\mathbf{X})\coloneqq[~{\boldsymbol{\phi}}(\mathbf{x}_{1})~~\cdots~~{\boldsymbol{\phi}}(\mathbf{x}_{m})~]\in\real^{n_{\phi}\times m}$. Importantly, the matrix $\mathbf{C}\in\real^{n_{y}\times n_{\phi}}$ can be computed once and reused repeatedly for online kernel evaluations. After constructing $\mathbf{C}$, the cost of evaluating a feature map kernel interpolant is therefore $\mathcal{O}(n_{\phi}n_{y})$, plus the expense of evaluating $\boldsymbol{\phi}$} once.
@@
${\boldsymbol{\phi}}(\mathbf{x})=\begin{bmatrix}\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\end{bmatrix}\in\real^{n_{x}+n_{x}^{2}},$ (2.13)
@@
$\mathbf{s}_{\mathbf{v}}^{\gamma}(\mathbf{x})=\mathbf{C}_{1}\mathbf{x}+\mathbf{C}_{2}[\mathbf{x}\otimes\mathbf{x}],\qquad\mathbf{C}=[~\mathbf{C}_{1}~~\mathbf{C}_{2}~]\in\real^{n_{y}\times(n_{x}+n_{x}^{2})}.$ (2.14)
@@ 2.3.3 Hybrid approach
where $c_{\phi},c_{\psi}\in\real_{>0}$ are positive weighting coefficients and $\boldsymbol{\phi}$} is chosen to keep $n_{\phi}$ from being too large. The resulting kernel interpolant then has the form
@@
where $\mathbf{A}_{1}\in\real^{n_{x}\times n_{x}}$, $\mathbf{A}_{2}\in\real^{n_{x}\times n_{x}^{2}}$, $\mathbf{A}_{3}\in\real^{n_{x}\times n_{x}^{3}}$, and $\mathbf{A}_{4}\in\real^{n_{x}\times n_{x}^{4}}$. One option is to fully capture the structure using a quartic feature map,
${\boldsymbol{\phi}}(\mathbf{x})=\begin{bmatrix}\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\otimes\mathbf{x}\\ \mathbf{x}\otimes\mathbf{x}\otimes\mathbf{x}\otimes\mathbf{x}\end{bmatrix}\in\real^{n_{x}+n_{x}^{2}+n_{x}^{3}+n_{x}^{4}}.$ (2.17)
@@
In some cases, in particular when using high-order polynomial feature maps, the kernel matrix $K(\mathbf{X},\mathbf{X})$ used for determining $\boldsymbol{\Omega}$} may be poorly conditioned. Increasing the regularization constant $\gamma$ can improve the conditioning of the system eq. 2.9b, but this can also degrade the accuracy of the resulting kernel interpolant. Applying a normalization to the inputs can help remedy the situation: for any injective ${\boldsymbol{\nu}}:\real^{n_{x}}\to\real^{n_{x}}$, if $K$ is positive definite, then the function $K_{\boldsymbol{\nu}}:\real^{n_{x}}\times\real^{n_{x}}\to\real$ defined by
@@
is also a positive-definite kernel function [50], and choosing $\boldsymbol{\nu}$} judiciously can improve the conditioning of $K_{\boldsymbol{\nu}}(\mathbf{X},\mathbf{X})$ compared to $K(\mathbf{X},\mathbf{X})$. A common choice is ${\boldsymbol{\nu}}(\mathbf{x})={\boldsymbol{\Sigma}}^{-1}(\mathbf{x}-\bar{\mathbf{x}})$, where ${\boldsymbol{\Sigma}}=\operatorname{diag}({\boldsymbol{\sigma}})\in\real^{n_{x}\times n_{x}}$ and $\bar{\mathbf{x}}\in\real^{n_{x}}$ with components
@@ 3.1 Generic projection-based reduced-order models
$\frac{\textrm{d}}{\textrm{d}t}\mathbf{q}(t)=\mathbf{f}(\mathbf{q}(t);\mathbf{u}(t)),\qquad\mathbf{q}(0)=\mathbf{q}_{0}({\boldsymbol{\mu}}),$ (3.1)
where $\mathbf{q}:[0,T]\to\real^{n_{q}}$ is the state, $\mathbf{u}:[0,T]\to\real^{n_{u}}$ is an input function, $\mathbf{f}:\real^{n_{q}}\times\real^{n_{u}}\to\real^{n_{q}}$ governs the state evolution, $\mathbf{q}_{0}({\boldsymbol{\mu}})\in\real^{n_{q}}$ is the initial condition parameterized by ${\boldsymbol{\mu}}\in{\cal D}\subset\real^{n_{\mu}}$, and $T>0$ is the final desired simulation time. Models of this form often arise from semi-discretizations of time-dependent partial differential equations (PDEs), in which case the large state dimension $n_{q}$ corresponds to the fidelity of the underlying mesh. We call eq. 3.1 the full-order model (FOM). For ease of exposition, we focus on autonomous systems without an input term $\mathbf{u}$; see Remark 4.2 for discussion on incorporating input terms.
@@
where $\mathbf{g}:\real^{r}\to\real^{n_{q}}$ and $\tilde{\mathbf{q}}:[0,T]\to\real^{r}$ is the reduced-order state, with $r\ll n_{q}$. The function $\mathbf{g}$ represents a decompression operation, mapping from reduced coordinates to the original high-dimensional space. We assume the existence of a corresponding compression map $\mathbf{h}:\real^{n_{q}}\to\real^{r}$, mapping high-dimensional states to reduced coordinates, such that $\mathbf{h}\circ\mathbf{g}:\real^{r}\to\real^{r}$ is the identity. Importantly, $(\mathbf{g}\circ\mathbf{h})^{2}=\mathbf{g}\circ(\mathbf{h}\circ\mathbf{g})\circ\mathbf{h}=\mathbf{g}\circ\mathbf{h}$, i.e., $\mathbf{g}\circ\mathbf{h}:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}}$ is a projection. The evolution for the reduced state $\tilde{\mathbf{q}}(t)$ is then given by
@@
in which $\mathbf{h}^{\prime}:\real^{n_{q}}\to\real^{n_{q}\times n_{q}}$ is the Jacobian of $\mathbf{h}$ and where the final step comes from inserting the approximation eq. 3.2 into the FOM eq. 3.1. The resulting system
@@ 3.2 Linear and quadratic dimension reduction
where $\bar{\mathbf{q}}\in\real^{n_{q}}$ is a fixed reference vector, $\mathbf{V}\in\real^{n_{q}\times r}$ has orthonormal columns, and $\mathbf{W}\in\real^{n_{q}\times r^{2}}$ satisfies $\mathbf{V}^{\mathsf{T}}\mathbf{W}=\bf 0$. This approximation defines an $r$-dimensional quadratic manifold embedded in $\real{}^{n_{q}}$. An appropriate compression map corresponding to eq. 3.5 is given by
@@
a system of $r\ll n_{q}$ ODEs defined by the function $\tilde{\mathbf{f}}:\real^{r}\to\real^{r}$.
@@
$\mathbf{q}_{k}^{(\ell)}\coloneqq\mathbf{q}(t_{k};{\boldsymbol{\mu}}_{\ell})\in\real^{n_{q}},\qquad\begin{aligned} \ell&=1,\ldots,M,\\ k&=0,1,\ldots,n_{t},\end{aligned}$ (3.9)
@@
First, if $\mathbf{W}=\bf 0$, the manifold defined by $\mathbf{g}$ has no curvature and reduces to an affine subspace (or a linear subspace if $\bar{\mathbf{q}}=\bf 0$) of $\real{}^{n_{q}}$. In this case, we select $\mathbf{V}$ using proper orthogonal decomposition (POD) [9, 22, 57]. Define
$\mathbf{Q}\coloneqq\begin{bmatrix}(\mathbf{q}_{0}^{(1)}-\bar{\mathbf{q}})&\cdots&(\mathbf{q}_{n_{t}}^{(1)}-\bar{\mathbf{q}})&(\mathbf{q}_{0}^{(2)}-\bar{\mathbf{q}})&\cdots&(\mathbf{q}_{n_{t}}^{(M)}-\bar{\mathbf{q}})\end{bmatrix}\in\real^{n_{q}\times M(n_{t}+1)},$ (3.11)
@@
The product $\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}$ contains redundant terms, i.e., $\tilde{q}_{i}\tilde{q}_{j}$ appears twice for each $i\neq j$, which means two columns of $\mathbf{W}\in\real^{r\times r^{2}}$ act on the same quadratic state interaction in the product $\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}$]. As a consequence, the learning problem eq. 3.12 has infinitely many solutions. In practice, this issue is avoided by replacing $\otimes$ in eq. 3.5 with a compressed Kronecker product $\tilde{\otimes}$, defined by
$\tilde{\mathbf{q}}\,\tilde{\otimes}\,\tilde{\mathbf{q}}\coloneqq\left[~\tilde{q}_{1}^{2}~~\tilde{q}_{1}\tilde{q}_{2}~~\tilde{q}_{2}^{2}~~\dots~~\tilde{q}_{r-1}\tilde{q}_{r}~~\tilde{q}_{r}^{2}~\right]^{\mathsf{T}}\in\real^{r(r+1)/2},$ (3.13)
which leads to a matrix $\tilde{\mathbf{W}}\in\real^{r\times r(r+1)/2}$ such that $\mathbf{W}[\tilde{\mathbf{q}}\otimes\tilde{\mathbf{q}}]=\tilde{\mathbf{W}}[\tilde{\mathbf{q}}\,\tilde{\otimes}\,\tilde{\mathbf{q}}$] for all $\tilde{\mathbf{q}}\in\real^{r}$. Then, if $\odot$ applies $\tilde{\otimes}$ column-wise, the optimization eq. 3.12 has a unique solution. Similar adjustments can be made for higher-order Kronecker products.
@@ 3.3 Intrusive reduced-order models for quadratic systems
where $\mathbf{A}\in\real^{n_{q}\times n_{q}}$, $\mathbf{H}\in\real^{n_{q}\times n_{q}^{2}}$. It is assumed that $\mathbf{H}$ is symmetric in the sense that $\mathbf{H}[\mathbf{q}\otimes\mathbf{p}]=\mathbf{H}[\mathbf{p}\otimes\mathbf{q}]$ for all $\mathbf{q},\mathbf{p}\in\real^{n_{q}}$. Models with quadratic structure arise from quadratic PDEs, but can also result from applying lifting transformations to models with other structure [31, 44]. With a linear state approximation ($\bar{\mathbf{q}}=\bf 0$ and $\mathbf{W}=\bf 0$), the ROM eq. 3.8 can be written as
@@
in which $\tilde{\mathbf{A}}=\mathbf{V}^{\mathsf{T}}\mathbf{A}\mathbf{V}\in\real^{r\times r}$ and $\tilde{\mathbf{H}}=\mathbf{V}^{\mathsf{T}}\mathbf{H}[\mathbf{V}\otimes\mathbf{V}]\in\real^{r\times r^{2}}$. Constructing eq. 3.15 is an intrusive process because $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{H}}$ depend explicitly on $\mathbf{A}$ and $\mathbf{H}$; however, we need not have access to $\mathbf{A}$ and $\mathbf{H}$ to observe that the quadratic structure is preserved.
@@
$\begin{aligned} \frac{\textrm{d}}{\textrm{d}t}\tilde{\mathbf{q}}(t)&=\tilde{\mathbf{A}}\tilde{\mathbf{q}}(t)+\tilde{\mathbf{H}}_{2}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{3}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)]+\tilde{\mathbf{H}}_{4}[\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)\otimes\tilde{\mathbf{q}}(t)],\\ \tilde{\mathbf{q}}(0)&=\mathbf{V}^{\mathsf{T}}\mathbf{q}_{0}({\boldsymbol{\mu}}),\end{aligned}$ (3.16)